OpenStudy (anonymous):

Medal Conchita missed 6 out of 65 questions on a medical terminology exam. What percent of the questions did she answer correctly? Round to the nearest tenth of a percent. ? %

sammixboo (sammixboo):

Okie dokie! Let's look at the formula. \(\large \rm \frac{part}{whole}~=~\frac{percent}{100}\) 6 is the part 65 is the whole \(\large \rm \frac{6}{65}~=~\frac{percent}{100}\) First take 6 and multiply it by 100 What is 6 x 100?

OpenStudy (anonymous):

600

sammixboo (sammixboo):

Good! Now we take 600 and divide it by 65 600/65 = your percent

sammixboo (sammixboo):

So what is 600/65? after you find that out we will round the answer to the nearest tenth

OpenStudy (anonymous):

9.23

OpenStudy (anonymous):

9.2307

sammixboo (sammixboo):

The tenths place is the first digit after the decimal point! 9.2307 - The 7 round the 0 up to 1, so now we have 9.231 9.231 - The one rounds the 3 to a 3, so now we have 9.23 9.23 - The 3 round the 2 to a 2, so now we have 9.2 9.2% is your answer!

OpenStudy (anonymous):

ohhhh i c now i understand it

jimthompson5910 (jim_thompson5910):

9.2% is the amount she missed subtract that from 100% to get the percent correct

OpenStudy (anonymous):

thank u sooooooooo much again

jimthompson5910 (jim_thompson5910):

you're on the right track though

OpenStudy (anonymous):

so the answer is 9.2%?

jimthompson5910 (jim_thompson5910):

you now need to evaluate 100% - 9.2%

OpenStudy (anonymous):

90.8

sammixboo (sammixboo):

Hmm 90% of 65 is way more than 6

jimthompson5910 (jim_thompson5910):

she missed 6 but she got 65 - 6 = 59 correct

OpenStudy (anonymous):

so whats the answer

jimthompson5910 (jim_thompson5910):

90.8% is the correct approximate answer

jimthompson5910 (jim_thompson5910):

6 wrong 59 correct 59/65 = 0.90769 ---> 90.769% which rounds to 90.8%

OpenStudy (anonymous):

ok thank you :)

sammixboo (sammixboo):

OH I misread the question. Thanks! @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're welcome